OpenStudy (jazzyfa30):

please help Exponential and Logarithmic Equations solve each equation. Give an exact solution. log49^x=(-1/2)

hartnn (hartnn):

is it \(\Large \log 49^x = \dfrac{-1}{2}\) ??

OpenStudy (jazzyfa30):

yes

hartnn (hartnn):

or \(\large \log_{49} x = -1/2 \) i think its prolly 2nd one

OpenStudy (jazzyfa30):

yes the second one sorry

hartnn (hartnn):

\(\large \log_ab=c \implies b= a^c\) \(\large \log_{49}x = -1/2 \implies ????\)

OpenStudy (jazzyfa30):

x=49^(-1/2)

hartnn (hartnn):

correct! now write 49 as 7^2 and try to simplify the exponent!

OpenStudy (jazzyfa30):

so it will 7^2^(-1/2)

hartnn (hartnn):

yes, notice that 2 will get cancelled what remains ?

OpenStudy (jazzyfa30):

how???

hartnn (hartnn):

 got that ?